OpenStudy (anonymous):

How do you complete an order of operation problem such as: 5 + (15-3) ????

OpenStudy (bahrom7893):

well depends on how you learned it.. but in this case it really doesnt matter: it is either 20 - 3 = 17 or 5+12 = 17

OpenStudy (anonymous):

Oh. Thank you bahrom7893!

OpenStudy (anonymous):

true order of operations says do the parenthesis first. so 5 + 12 = 17

OpenStudy (bahrom7893):

densmath it doesn't matter